Jump to content

Recommended Posts

We are aware of the issue with the NOC, Bill and our Data Centre Manager are checking it out. We will update you soon.

Kindly accept our apologies for the inconveniences caused.

Thanks for the quick notice Carl. Triggered our sensors 3 minutes ago ;).

Link to post
Share on other sites

We have been monitoring the issue, everything is back to normal after the swap earlier. If you are still facing issues, please contact help desk using the link on the top menu or on my signature, we will gladly check it out.

 

Closing thread.

Link to post
Share on other sites
Guest
This topic is now closed to further replies.
×
×
  • Create New...